OpenStudy (anonymous):

When 8 is subtracted from Ollie’s age, the result is the same as subtracting 16 from 3 times his age. How old is Ollie? ?=years I think it is 13

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

Wait give me a second.

OpenStudy (anonymous):

deffinetly not that

OpenStudy (anonymous):

Ik that now

OpenStudy (anonymous):

i got it just one second

OpenStudy (anonymous):

Let Ollies age be x. We know that x-8 is equal to 3x-16.

OpenStudy (anonymous):

solving for x, we see that 2x=8

OpenStudy (anonymous):

I think it is 4.

OpenStudy (anonymous):

Therefore, x is 4

OpenStudy (anonymous):

Cole1117 is correct

OpenStudy (anonymous):

it is 4

OpenStudy (anonymous):

Is it 4 or some where around that area/................. thx guys

OpenStudy (anonymous):

little turds had to come in and take over when someone was already helping lol

OpenStudy (anonymous):

LolZz

OpenStudy (anonymous):

4-8=-4 4x3-16=-4

OpenStudy (anonymous):

Jeff, the correct answer is 4. 4-8= -4. 4*3=12-16=-4